OpenStudy (anonymous):

graph y=4cos(2x-π). be sure to label your axis. state the amplitude, period, phase shift and vertical shift

OpenStudy (jhannybean):

Your function is in the form : \(F(x) = Ay(Bt-C)\) - amplitude : \(A\) - period: \(\dfrac{2\pi}{|B|}\) - phase shift : \(\dfrac{C}{B}\) - vertical shift: none

OpenStudy (jhannybean):

http://www.purplemath.com/modules/grphtrig.htm :)